# Cold-Start Recommendation with Knowledge-Guided Retrieval-Augmented Generation

### v1

@@ Abstract.
conference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems have come to play a pivotal role in various domains, such as e-commerce, streaming services, and online advertising, by helping users navigate an overwhelming amount of content (Sarwar et al., 2001; Koren et al., 2009). However, despite decades of progress in collaborative filtering (CF) and neural recommendation models, these methods still struggle in cold-start scenarios, where new users or items have little to no historical interaction data (Schein et al., 2002; Park and Chu, 2009; Gope and Jain, 2017; Zhang et al., 2025). Because CF methods rely on user–item interaction data, their predictive power diminishes significantly when such data are sparse or unavailable. Content-based (CB) approaches help alleviate data sparsity by leveraging side information such as item attributes and user profiles. However, their effectiveness is limited by the completeness and reliability of the available metadata (Lam et al., 2008; Volkovs et al., 2017b; Gu et al., 2016). Hybrid models that combine collaborative filtering with content-based methods (CF + CB) can partially mitigate the cold-start problem, but they still require some level of interaction history to perform effectively (Volkovs et al., 2017a; Zhang and Wang, 2015). Consequently, cold-start remains a major obstacle to delivering accurate recommendations in dynamic environments where new users and items are continuously introduced.
@@
Recent advances in large language models (LLMs) have opened a promising zero-shot avenue: framing recommendation as a natural-language generation task and prompting the LLM to infer relevant items directly (Kojima et al., 2022; He et al., 2023; Hou et al., 2024). LLMs possess rich world knowledge acquired during pretraining, enabling them to make reasonable inferences even with minimal domain-specific data. Since the cold-start problem requires inferring user preferences with no historical interaction data, it can be viewed as a subset of the zero-shot setting (Li et al., 2019). Due to this connection, several studies have explored LLM-based approaches to address the cold-start challenge. Recent studies show that prompting LLMs to rank candidate items purely from textual cues—sometimes augmented with lightweight keyword extraction or retrieval steps—can deliver strong cold-start performance without any domain-specific fine-tuning (Hou et al., 2024; Liang et al., 2025; Kieu et al., 2024). Together, these works demonstrate that zero-shot LLMs can rival or surpass specialized recommenders in data-sparse settings, inspiring continued exploration of retrieval-augmented and hierarchy-aware prompting strategies. A different approach, inspired by tree-based reasoning, organizes the candidate items into a hierarchical structure and prompts the LLM to iteratively refine the recommendations by eliminating less relevant options at each step (Zhang et al., 2024). These approaches demonstrate the potential of LLMs to mitigate the cold-start problem by enabling zero-shot recommendation, offering a promising direction for improving personalization when user-item interaction data is limited.
However, naive prompt-based LLM recommenders suffer from three critical limitations (Huang et al., 2025b). (i) Hallucination: They can hallucinate facts or item attributes that are not present in the underlying catalog, resulting in invalid or out-of-domain suggestions (Jiang et al., 2025; Islam and Faruk, 2025). To constrain generation strictly to in-domain content, prior approaches often compile a candidate list of admissible items and prepend it to the prompt. However, the LLM’s finite context window necessitates aggressive truncation of this list, crowding out essential user and item cues and ultimately degrading recommendation quality. (ii) Outdated knowledge : Their implicit knowledge may not reflect the most recent or niche inventory, especially in rapidly changing domains where new items appear continuously (Huang et al., 2025a). (iii) Generation instability: The stochastic decoding processes that underpin modern LLMs make their recommendations inherently non-deterministic: even minor prompt perturbations or sampling-temperature changes can produce markedly different ranked lists, hampering reproducibility, complicating offline evaluation, and undermining user trust in production (Holtzman et al., 2020). To address these limitations, recent work has explored retrieval-augmented generation (RAG) pipelines, which ground an LLM’s output in external evidence by retrieving documents relevant to the user’s query (Lewis et al., 2020; Peng et al., 2024). Existing RAG recommenders typically rely on flat text corpora and lack an explicit mechanism for multi-hop reasoning over structured relationships that link items together (e.g., shared genres, themes, or creators)-relationships that are often crucial for identifying semantically similar cold items (Kieu et al., 2024).
In this paper we introduce ColdRAG, a retrieval-augmented generation framework that unifies knowledge graph reasoning with LLM recommendation to address the item cold-start challenge without any task-specific fine-tuning. To effectively recommend previously unseen items, ColdRAG proceeds through four sequential stages. (i) Item profile generation: It converts available structured fields for each item-such as titles, attributes, and reviews—into a coherent natural-language profile, enriching sparse metadata with the LLM’s background knowledge. (ii) Knowledge graph construction: It extracts typed entities and relations from these profiles to form a domain-specific knowledge graph that explicitly encodes cross-item semantics. (iii) Multi-hop reasoning-based retrieval: Given a user’s interaction history as a query, the system traverses this graph under LLM guidance to retrieve candidate items along with supporting evidence, and finally prompts the LLM to rank them. (iv) Recommendation with retrieval-augmented generation: The LLM ranks the graph-retrieved candidates using their associated supporting facts and composes a concise, source-cited recommendation list. By grounding generation in evidence, ColdRAG mitigates hallucinations, supports multi-hop reasoning, and seamlessly incorporates new items.
We highlight the advantages of ColdRAG through a concise comparison with other recommender system categories in Table 1. In this table, the Cold-start column indicates the capability to recommend previously unseen items without relying on historical user–item interactions. Grounding refers to the use of external knowledge sources or retrieved information to support recommendations. Zero-shot denotes whether the model can be applied directly without any task-specific fine-tuning. The Multi-hop column reflects the model’s capacity to perform reasoning through traversal over knowledge graphs. In summary, our contributions are threefold:
  • We identify limitations of existing zero-shot LLM recommenders in cold-start scenarios and motivate the need for structured reasoning over explicit knowledge.
@@
  • We propose ColdRAG that integrates retrieval-augmented generation with dynamically constructed knowledge-graph reasoning for recommendation—without requiring any task-specific training.
  • We achieve state-of-the-art zero-shot performance on multiple benchmarks and conduct detailed analyses underscoring the generation stability advantages of retrieval grounded in knowledge graphs.
@@ 2.1. Cold-Start Recommendation
The cold-start problem is a well-recognized challenge in recommendation systems, occurring when new users or items lack sufficient historical interaction data. Early approaches relied on content-based or hybrid techniques that leverage auxiliary information—such as item descriptions or user attributes—to compensate for missing collaborative signals. Representative examples include collaborative deep learning, which jointly embeds textual content and interaction data (Wang et al., 2015), and meta-learning frameworks that treat cold-start as a few-shot learning problem (Lee et al., 2019). These methods illustrate a broader trend toward transferring knowledge from side information or related tasks to warm up recommendations for new entities.
More recent efforts focus on enhancing neural recommender models with mechanisms specifically designed for cold-start scenarios. Some methods introduce stochastic training strategies: DropoutNet, for instance, explicitly simulates cold-start by randomly dropping out warm-item embeddings during training to learn robust representations for new items (Volkovs et al., 2017a). Other approaches add regularization or alignment objectives to ensure consistency between representations of warm and cold items. Contrastive learning has proven effective in this regard—CLCRec uses contrastive objectives to preserve collaborative signals in content-derived embeddings, significantly improving cold-start item recommendations (Wei et al., 2021). Distributional robustness offers another promising direction, with TDRO applying distributionally robust optimization to account for temporal shifts in item features and thereby improve generalization to unseen items (Lin et al., 2024). Meanwhile, general sequential recommenders like SASRec (Kang and McAuley, 2018) that rely on item IDs tend to struggle on unseen items, a limitation addressed by methods such as UniSRec (Hou et al., 2022) which replaces IDs with textual embeddings to enable cross-domain transfer. Still, these models either require dedicated cold-start modules, extensive side information, or retraining when new items emerge. In contrast, our proposed ColdRAG framework avoids specialized retraining by leveraging a LLM with retrieval augmentation, enabling zero-shot recommendations for new items through generative reasoning over relevant item content.
@@ 2.2. LLM-based Recommender Systems
Viewing recommendation through the lens of natural-language processing has recently attracted significant attention. Early efforts such as P5 treat diverse recommendation tasks as a unified text-to-text problem, fine-tuning a T5 backbone to produce recommendations, explanations, and other outputs with a single model (Geng et al., 2022). Subsequent studies exploit the semantic richness of pre-trained language encoders to embed item descriptions, reviews, or user profiles, improving data efficiency and cross-domain transfer.
With the advent of powerful LLMs, researchers have explored prompt-based, zero-shot recommenders. LLMRank casts ranking as instruction following: given a user’s history and candidate descriptions, an LLM ranks items without additional training (Hou et al., 2024). PromptRec shows that carefully designed prompts enable even relatively small language models to handle extreme cold-start settings (Wu et al., 2024). TaxRec augments prompts with taxonomy tokens so the LLM can better navigate hierarchical item spaces (Liang et al., 2025). KALM4Rec couples a graph-based keyword retrieval module with prompt-based LLM re-ranking for cold-start user recommendation without any additional fine-tuning(Kieu et al., 2024). While these models demonstrate impressive zero-shot ability, they rely solely on the prompt’s immediate text and thus can hallucinate or miss domain-specific details. Our ColdRAG differs by explicitly retrieving and supplying relevant item knowledge—structured as a dynamically built graph—to the LLM, grounding its reasoning and improving accuracy for newly introduced items.
@@ 3. Proposed Method
We propose a Retrieval-Augmented Generation (RAG) framework tailored for item cold-start recommendation, called ColdRAG. The key idea is to equip a large language model (LLM) with external knowledge, represented as a graph, to guide multi-hop reasoning beyond its internal knowledge from pre-trained data. This combination enables both semantic reasoning and evidence-grounded generation for recommending cold-start items. Rather than relying solely on internal LLM knowledge or static prompts, our approach dynamically builds a knowledge graph from item metadata and engages the LLM in structured reasoning and response composition based on retrieved information. Figure 1 illustrates the overall pipeline, which consists of four major stages: (1) item profile generation, (2) knowledge graph construction, (3) multi-hop reasoning-based retrieval, and (4) recommendation with retrieval-augmented generation.
@@ 3.2. Item Profile Generation
This item profile serves as the foundation for constructing the knowledge graph. Its structured narrative enables the accurate extraction of entities and relations, which form the nodes and edges of the graph. By abstracting away noise and standardizing item-level semantics, the framework ensures that the resulting graph is well-suited for multi-hop reasoning—particularly in item cold-start scenario.
@@ 3.5. Recommendation with Retrieval-augmented Generation
Table 4. Comparison of Recall@10 and NDCG@10 across three domains: Games, Toys, and Office. The final row shows the improvement of our best-performing model compared to the best baseline per domain. Our proposed models are highlighted in gray. For each column, the best result is highlighted in bold, and the second-best is underlined. All results for the zero-shot models are averaged over 5 runs.
@@ 4. Experiments
We conduct comprehensive experiments to evaluate the effectiveness of ColdRAG in addressing the item cold-start recommendation challenge. To guide our analysis, we organize our evaluation around the following five research questions:
@@ 4.1.1. Datasets
To simulate the cold-start setting, we identify the bottom 10% least frequently interacted items in each dataset as cold items. We follow a sequential recommendation setting, treating each user’s interaction history as a sequence. Specifically, we adopt the widely used leave-one-out evaluation (Sun et al., 2019; Hou et al., 2022), where the last item in each user sequence is held out as the target for prediction. From each dataset, we sample 500 user sequences that end with a cold item. For each sequence, the first $n{-}1$ items are used as input to the model, and the final $n$th item is used as the test target.
For full-train baselines, we use the 500 cold-item user sequences as the test set and treat all remaining user sequences in the dataset as the training set. This design allows us to fairly compare zero-shot and fully supervised (full-train) methods under consistent evaluation protocols, while ensuring that cold items are properly isolated at inference phase. The statistics of the datasets after preprocessing are summarized in Table 2. The knowledge graph statistics constructed from the item profiles are reported in Table 3.
@@ 4.1.2. Baselines
  • UniSRec (Hou et al., 2022): A universal sequence representation learning model that leverages contrastive learning and item text encoders to model user behavior. We fine-tune the pre-trained UniSRec model on our domain-specific training data. This baseline serves as a strong representation-based recommender for comparison.
@@ 4.1.3. Evaluation Metrics
To evaluate the performance of the recommendation, we use two metrics widely adopted in cold-start recommendation tasks: Recall@$k$ and NDCG@$k$, where $k=10$. Recall@10 measures whether the ground-truth item appears within the top-10 predicted items, reflecting the model’s ability to retrieve relevant results. NDCG@10 (Normalized Discounted Cumulative Gain) additionally considers the position of the ground-truth item in the ranked list, assigning higher scores to correct items ranked closer to the top.
@@ 4.2.1. Overall Performance (RQ1)
Table 4 presents the performance comparison across all baseline models and our proposed method on three domains: Games, Toys, and Office. As shown, ColdRAG consistently achieves the highest performance across all datasets and metrics, demonstrating its effectiveness in addressing the item cold-start challenge. The model outperforms both full-train and zero-shot baselines, achieving an average improvement of +31.31% Recall@10 on Games, +21.11% on Toys, and +55.81% on Office compared to the best-performing baseline in each domain.
For the full-train baselines, TDRO achieves the highest performance by leveraging temporally and distributionally robust optimization to address data distribution shifts. However, it is still outperformed by ColdRAG across all domains. For instance, in the Office domain, ColdRAG achieves a Recall@10 score that is substantially higher than TDRO, indicating its superior ability to recommend cold-start items under sparse conditions. These results suggest that even advanced fully supervised models face limitations in capturing item-level semantics in cold-start setting, whereas ColdRAG’s retrieval-augmented, knowledge-grounded design enables more effective and adaptive recommendation.
Among the zero-shot baselines, the plain LLM and LLMRank variants (LLMRank (S), (R), and (I)) demonstrate varying degrees of effectiveness depending on how they prompt the LLM, yet all rely on flat prompts and pre-defined candidate lists. This limitation hinders their ability to adaptively explore semantic relationships beyond what is explicitly given in the prompt. KALM4Rec performs better by leveraging keyword-based retrieval to enrich prompts with relevant context. However, ColdRAG surpasses all these approaches by integrating structured, knowledge-grounded reasoning through multi-hop traversal over a dynamically constructed knowledge graph. This allows it to retrieve richer and more semantically aligned item context, leading to superior recommendation quality.
In summary, the strong and consistent performance of ColdRAG across both full-train and zero-shot baselines highlights the effectiveness of integrating structured external knowledge and graph-based reasoning into LLM-based recommendation. By combining graph traversal with language understanding, ColdRAG achieves not only higher accuracy but also greater adaptability in cold-start scenarios where data is limited. These results affirm the strength of ColdRAG’s design and its promise as a robust solution for zero-shot recommendation in item cold-start scenario.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems underpin e-commerce, streaming, and advertising (Sarwar et al., 2001; Koren et al., 2009), yet performance in the cold-start setting—where new users or items have little to no interaction history—remains fragile (Gope and Jain, 2017; Zhang et al., 2025). Collaborative filtering (CF) depends on dense user–item signals and degrades sharply under sparsity. Content-based (CB) approaches leverage side information (e.g., item attributes, user profiles) but are constrained by incomplete or noisy metadata (Lam et al., 2008; Volkovs et al., 2017b; Gu et al., 2016). Hybrid models (CF+CB) partly alleviate these issues, yet still require some interactions to anchor predictions (Volkovs et al., 2017a; Zhang and Wang, 2015). Consequently, robust cold-start recommendation remains a central obstacle as catalogs and user populations evolve continuously.
Recent advances in large language models (LLMs) have opened a zero-shot avenue to recommendation by casting it as natural-language generation and prompting the model to directly infer relevant items (Kojima et al., 2022; Hou et al., 2024; Li et al., 2019; Xu et al., 2025). Because the cold-start setting is inherently zero-shot, a growing body of work explores LLM-based strategies for new users and items, leveraging pretrained world knowledge with little or no domain-specific data (Li et al., 2019; Hou et al., 2024). In particular, prompting LLMs to score or rank candidates from textual cues—optionally aided by lightweight keyword extraction or retrieval—achieves competitive cold-start performance without fine-tuning (Hou et al., 2024; Liang et al., 2025; Kieu et al., 2024). Complementary approaches organize candidates hierarchically and iteratively prune them with tree-style reasoning to sharpen recommendations (Zhang et al., 2024). Despite this promise, prompt-only LLM recommenders remain vulnerable to hallucination, stale or niche inventory coverage gaps, and generation instability, underscoring the need for retrieval augmentation and robustness techniques (Huang et al., 2025b; Jiang et al., 2025; Islam and Faruk, 2025; Huang et al., 2025a; Holtzman et al., 2020; Yang et al., 2024).
@@
Retrieval-augmented generation (RAG) grounds outputs in external evidence (Lewis et al., 2020; Peng et al., 2024), but common variants misalign with cold-start setting. Similarity-embedding RAG flattens structure and returns top-$k$ textually similar items that may be semantically adjacent yet decision-irrelevant. Naive GraphRAG pipelines inject structure but either rely on global community summaries—useful for topic-level queries yet too coarse for item-level personalization—or perform $k$-hop expansion that sweeps in all neighbors within $k$ hops (Xu et al., 2024; Edge et al., 2024; Guo et al., 2024), inducing exponential neighborhood growth that quickly exhausts context and dilutes the few relations that justify a cold-start choice (Xu et al., 2018; Topping et al., 2021; Lee et al., 2025). In cold-start, what matters is not neighborhood size but which relations and paths (e.g., shared genres, themes, creators) carry explanatory power (Kieu et al., 2024).
We target an LLM-augmented, KG-based recommender that is robust to item cold-start settings. To this end, we pursue three design goals: Knowledge grounding & generation stability—constrain outputs to KG-retrieved candidates with explicit evidence, reducing hallucinations and stabilizing outputs; Multi-hop Reasoning-based Retrieval—prioritize semantically meaningful relation paths over bulk neighborhoods to form a compact, high-quality candidate set; and Cold-start adaptability with explainability—integrate new items into the KG for immediate recommendations, accompanied by concise, path-based rationales.
In this work, we propose ColdRAG that casts retrieval as goal-directed path selection: during traversal, relations are scored for how much they advance the recommendation objective, and branches that add little explanatory value are pruned early. This converts KG structure from a passive index into an active reasoning space that builds compact evidence chains under a fixed token budget. ColdRAG proceeds in four stages: (i) Item profile generation converts available fields (titles, attributes, reviews) into coherent natural-language profiles; (ii) Knowledge graph construction extracts typed entities and relations to form a domain-specific KG; (iii) Multi-hop reasoning-based retrieval scores relation paths and prunes unrelated branches to retrieve a compact candidate set with supporting evidence; and (iv) Recommendation with retrieval-augmented generation ranks candidates using associated supporting facts to produce a concise, source-cited list. By grounding generation in explicit evidence and constraining traversal to high-yield paths, ColdRAG mitigates hallucination, incorporates new items seamlessly, and improves stability—without incurring exponential neighborhood expansion.
We highlight the advantages of ColdRAG in Table 1, which shows representative categories (CF/CB/Hybrid, prompt-only LLMs, and RAG variants) along four axes: Cold-start, Grounding, Zero-shot, and Multi-hop. In summary, our contributions are threefold:
  • We diagnose why prompt-only LLM recommenders and conventional RAG (similarity- and $k$-hop–based) underperform in cold-start, motivating structured reasoning over explicit knowledge.
@@
  • We propose ColdRAG, integrating retrieval-augmented generation with dynamically constructed KG reasoning and path-selective, budget-aware retrieval that returns evidence chains without task-specific training.
  • We demonstrate state-of-the-art zero-shot results on multiple benchmarks and analyze the stability gains attributable to path-selective retrieval and evidence-grounded generation.
@@ 2.1. Cold-Start Recommendation
The cold-start problem (Zhang et al., 2025) is a well-recognized challenge in recommendation systems, occurring when new users or items lack sufficient historical interaction data, prompting early content-based or hybrid techniques that leverage auxiliary information—such as item descriptions or user attributes—to compensate for missing collaborative signals. Representative approaches include collaborative deep learning, which jointly embeds textual content and interaction data (Wang et al., 2015), and meta-learning frameworks that treat cold-start as a few-shot learning problem (Lee et al., 2019), reflecting a broader trend toward transferring knowledge from side information or related tasks. More recent efforts add explicit cold-start mechanisms: DropoutNet simulates cold-start by randomly dropping out warm-item embeddings during training to learn robust representations (Volkovs et al., 2017a), while contrastive learning methods like CLCRec preserve collaborative signals in content-derived embeddings to improve new-item recommendations (Wei et al., 2021). Distributional robustness offers another promising direction, with TDRO applying distributionally robust optimization to account for temporal shifts in item features and thereby enhance generalization (Lin et al., 2024). General sequential recommenders like SASRec (Kang and McAuley, 2018) that rely on item IDs tend to struggle on unseen items, a limitation addressed by UniSRec (Hou et al., 2022) which replaces IDs with textual embeddings to enable cross-domain transfer. Still, these models either require dedicated cold-start modules, extensive side information, or retraining when new items emerge, whereas our proposed ColdRAG framework avoids finetuning by leveraging an LLM with retrieval augmentation to enable zero-shot recommendations through generative reasoning.
@@ 2.2. LLM-based Recommender Systems
Early efforts such as P5 treat diverse recommendation tasks as a unified text-to-text problem, fine-tuning a T5 backbone to produce recommendations, explanations, and other outputs with a single model (Geng et al., 2022). Subsequent studies exploit the semantic richness of pre-trained language encoders to embed item descriptions, reviews, or user profiles, improving data efficiency and cross-domain transfer. With the advent of powerful LLMs, researchers have explored prompt-based, zero-shot recommenders. LLMRank casts ranking as instruction following: given a user’s history and candidate descriptions, an LLM ranks items without additional training (Hou et al., 2024). PromptRec shows that carefully designed prompts enable even relatively small language models to handle extreme cold-start settings (Wu et al., 2024). TaxRec augments prompts with taxonomy tokens so the LLM can better navigate hierarchical item spaces (Liang et al., 2025). KALM4Rec couples a graph-based keyword retrieval module with prompt-based LLM re-ranking for cold-start user recommendation without fine-tuning (Kieu et al., 2024). While these models demonstrate impressive zero-shot ability, they rely solely on the prompt’s immediate text and thus can hallucinate or miss domain-specific details. Our ColdRAG differs by explicitly retrieving and supplying relevant item knowledge—structured as a dynamically built graph—to the LLM, grounding its reasoning and improving accuracy for newly introduced items.
@@ 3. Proposed Method
We propose a Retrieval-Augmented Generation (RAG) framework tailored for item cold-start recommendation, called ColdRAG. The key idea is to equip a large language model (LLM) with external knowledge, represented as a graph, to guide multi-hop reasoning beyond its internal knowledge from pre-trained data. This combination enables both semantic reasoning and evidence-grounded generation for recommending cold-start items. Rather than relying solely on internal LLM knowledge or static prompts, our approach dynamically builds a knowledge graph from item metadata and engages the LLM in structured reasoning and response composition based on retrieved information. Figure 1 illustrates the pipeline with four major stages: (1) Item Profile Generation, (2) Knowledge Graph Construction, (3) Multi-hop Reasoning-based Retrieval, and (4) Recommendation with Retrieval-augmented Generation.
@@ 3.2. Item Profile Generation
This item profile serves as the foundation for constructing the knowledge graph. Its structured narrative enables the accurate extraction of entities and relations, which form the nodes and edges of the graph. By abstracting noise and standardizing item-level semantics, the framework ensures that the resulting graph is well-suited for multi-hop reasoning—particularly in item cold-start scenario.
@@ 3.5. Recommendation with Retrieval-augmented Generation
Table 4. Comparison of Recall@10 and NDCG@10 across three datasets: Games, Toys, and Office. Our proposed models are highlighted in gray. For each column, the best result is highlighted in bold, and the second-best is underlined. The final row shows the improvement of our best-performing model compared to the best baseline per dataset. All results are averaged over 5 runs.
@@ 4. Experiments
We conduct comprehensive experiments to evaluate the effectiveness of ColdRAG in item cold-start recommendation. To guide our analysis, we organize our evaluation around the following five research questions:
@@ 4.1.1. Datasets
To simulate the cold-start setting, we identify the bottom 10% least frequently interacted items in each dataset as cold items. We follow a sequential recommendation setting, treating each user’s interaction history as a sequence. Specifically, we adopt the widely used leave-one-out evaluation (Sun et al., 2019; Hou et al., 2022), where the last item in each user sequence is held out as the target for prediction. From each dataset, we sample 500 user sequences that end with a cold item. For each sequence, the first n1n{-}1 items are used as input to the model, and the final $n$th item is used as the test target.
For full-train baselines, we use the 500 cold-item user sequences as the test set and treat all remaining user sequences in the dataset as the training set. This design allows us to fairly compare zero-shot and fully supervised (full-train) methods under consistent evaluation protocols, while ensuring that cold items are properly isolated at inference phase. The statistics of the datasets after preprocessing are summarized in Table 2. The knowledge graph statistics constructed from the item profiles are reported in Table 3, with a detailed breakdown of the entity and edge type distributions provided in Figure 5.
@@ 4.1.2. Baselines
  • UniSRec (Hou et al., 2022): A universal sequence representation learning model that leverages contrastive learning and item text encoders to model user behavior. We fine-tune the pre-trained UniSRec model on domain-specific data. This baseline serves as a strong representation based recommender for comparison.
@@ 4.1.3. Evaluation Metrics
To evaluate the performance of the recommendation, we use two metrics widely adopted in cold-start recommendation tasks: Recall@$k$ and NDCG@$k$. Following prior works (Hou et al., 2022; Wei et al., 2021; Liang et al., 2025), we report results at $k=10$.
@@ 4.2.1. Overall Performance (RQ1)
Table 4 presents the performance comparison across all baseline models and our proposed method on three datasets: Games, Toys, and Office. As shown, ColdRAG consistently achieves the highest performance across all datasets and metrics, demonstrating its effectiveness in addressing the item cold-start challenge. The model outperforms both full-train and zero-shot baselines, achieving an average improvement of +39.62% Recall@10 on Games, +26.76% on Toys, and +104.76% on Office compared to the best-performing baseline in each domain. Notably, the improvements in Recall are consistently more pronounced than those in NDCG. This suggests that ColdRAG’s primary strength lies in its multi-hop reasoning-based retrieval, which excels at identifying and including the correct item within the top candidates.
For the full-train baselines, TDRO achieves overall high performance by leveraging temporally and distributionally robust optimization to address data distribution shifts. However, it is still outperformed by ColdRAG across all domains. For instance, in the Office domain, ColdRAG achieves a Recall@10 score that is substantially higher than TDRO, indicating its superior ability to recommend cold-start items under sparse conditions. These results suggest that even advanced fully supervised models face limitations in capturing item-level semantics in cold-start setting, whereas ColdRAG’s retrieval-augmented, knowledge-grounded design enables more effective and adaptive recommendation.
Among the zero-shot baselines, the plain LLM and LLMRank variants (LLMRank (S), (R), and (I)) demonstrate varying degrees of effectiveness depending on how they prompt the LLM, yet all rely on prompts and pre-defined candidate lists. This limitation hinders their ability to adaptively explore semantic relationships beyond what is explicitly given in the prompt. KALM4Rec performs better by leveraging keyword-based retrieval to enrich prompts with relevant context. However, ColdRAG surpasses all these approaches by integrating structured, knowledge-grounded reasoning through multi-hop traversal over a dynamically constructed knowledge graph. This allows it to retrieve richer and more semantically aligned item context, leading to superior recommendation quality.
In summary, the strong and consistent performance of ColdRAG across both full-train and zero-shot baselines highlights the effectiveness of integrating structured external knowledge and graph-based reasoning into LLM-based recommendation. These results affirm the strength of ColdRAG’s design and its promise as a robust solution for zero-shot recommendation in item cold-start scenario.